OpenStudy (anonymous):

Describe the graphs of the functions f(x) = 2x + 1 and g(x) = –2x + 1

OpenStudy (anonymous):

is this correct: f(x)=2^x 1= y=2^x 1 f(x)=-2^x 1 I put in values in for x and y. y=2(0) 1=(0,1) y=2(-2) 1=(-2,-3) -------------------------------------------------------- g(x)=-2(x)+1=y=-2(x)+1 y=-2(0)+1=y=1(0,1) y=-2(-2)+1=y=5(-2,5)